# Similarity-Distance-Magnitude Activations

### v2

@@ 5 Results
The results for the in-distribution $\operatorname{\text{{Sentiment}}}$ dataset appear in Table 1. Even on this in-distribution dataset, the estimators over the underlying LMs without adaptor layers exhibit modest over-confidence, which is reflected in conditional accuracies that fall below the expected $\alpha$. The estimators over the adaptor layers and the $\operatorname{\text{{sdm}}}$ activation all obtain the desired conditional accuracies, with the class-wise accuracies of the models themselves $\geq\alpha$, with differences arising in the proportion of admitted points. Here and elsewhere, $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ tends to be overly conservative in rejecting points. To be expected, the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ estimator tends to be more conservative than simply thresholding the $\operatorname{\text{{sdm}}}$ activation on the in-distribution data, but the latter lacks the assurances on the class-conditional accuracy obtained by the constraints on the $\operatorname{\text{{high-reliability}}}$ region. In practice, this behavior can be used as a basis to triage selective classifications: For example, documents in the $\operatorname{\text{{high-reliability}}}$ region might be treated as automated, or semi-automated, predictions in the decision pipeline, whereas other documents might be triaged by $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})_{\hat{y}}$ for calling more resource-intensive LM tools, or human adjudication. Importantly, as we discuss below with the co-variate-shifted and out-of-distribution datasets, just using the $\operatorname{softmax}$ or the other estimators does not provide a reliable substrate for basing such conditional branching decisions.
@@
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.99 0.43 0.99 0.38 0.99 0.43 0.99 0.38 0.99 0.81
@@
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.98 0.48 0.98 0.43 0.98 0.47 0.98 0.43 0.98 0.90
@@
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. 0.01 0.98 0.01 0.98 0.01 1. 0.01 0.99 0.02
@@
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.98 0.05 0.96 0.04 0.97 0.05 0.98 0.04 0.97 0.08
@@
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. 0.01 0.97 0.14 0.75 0.02 1. 0.14 0.97 0.16
@@
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.88 0.10 0.95 0.18 0.91 0.09 0.93 0.18 0.93 0.27
@@
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}},\alpha=0.94$ 0.85 0.11 0.95 0.18 0.92 0.10 0.91 0.19 0.91 0.29
@@
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.99 0.42 0.99 0.39 0.99 0.42 0.99 0.39 0.99 0.81
@@
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.99 0.47 0.98 0.43 0.98 0.47 0.98 0.43 0.98 0.90
@@
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.98 0.33 0.99 0.27 0.99 0.33 0.98 0.27 0.98 0.60
@@
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.97 0.41 0.95 0.32 0.96 0.41 0.95 0.32 0.96 0.73
@@
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}},\alpha=0.94$ 0.96 0.42 0.95 0.33 0.96 0.42 0.95 0.32 0.96 0.74

### v3

@@ 5 Results
Representative results are provided in Table 1, with additional results in the Appendix. Even on the in-distribution $\operatorname{\text{{Sentiment}}}$ dataset, the estimators over the underlying LMs without adaptor layers exhibit over-confidence, which is reflected in conditional accuracies that fall below the expected $\alpha$. The estimators over the adaptor layers all obtain the desired conditional accuracies, with the class-wise accuracies of the models themselves $\geq\alpha$ (see the $\operatorname{\text{{no-reject}}}$ rows in Table 2), with differences arising in the proportion of admitted points. Here and elsewhere, $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ tends to be overly conservative in rejecting points. To be expected, the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ estimator tends to be more conservative than simply thresholding the $\operatorname{\text{{sdm}}}$ activation at $\alpha$ ($\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$), but the latter lacks the assurances on the class-conditional accuracy obtained by the constraints on the $\operatorname{\text{{high-reliability}}}$ region. In practice, this behavior can be used as a basis to triage selective classifications: Documents in the $\operatorname{\text{{high-reliability}}}$ region might be treated as automated, or semi-automated, predictions in the decision pipeline, whereas other documents might be triaged by $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})_{\hat{y}}$ for calling more resource-intensive LM tools, or human adjudication. Importantly, as we discuss below with the co-variate-shifted and out-of-distribution datasets, the non-$\operatorname{\text{{sdm}}}$-based estimators provide a less reliable substrate for basing such conditional branching decisions.
Class-conditional Prediction-conditional Marginal
$y=0$ $y=1$ $\hat{y}=0$ $\hat{y}=1$ $y\in\{0,1\}$
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+DiscVBLLMLP}}}$ $\operatorname{\text{{vbll}}}$ 0.99 0.42 1.00 0.40 1.00 0.42 0.99 0.40 0.99 0.82
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 0.99 0.43 0.99 0.38 0.99 0.43 0.99 0.38 0.99 0.81
[…]
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+DiscVBLLMLP}}}$ $\operatorname{\text{{vbll}}}$ 0.99 0.44 0.99 0.43 0.99 0.44 0.99 0.43 0.99 0.87
[…]
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+DiscVBLLMLP}}}$ $\operatorname{\text{{vbll}}}$ 0.89 0.03 0.96 0.04 0.93 0.02 0.94 0.05 0.94 0.07
[…]
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+DiscVBLLMLP}}}$ $\operatorname{\text{{vbll}}}$ 0.94 0.01 0.97 0.11 0.79 0.02 0.99 0.10 0.96 0.12
[…]
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+DiscVBLLMLP}}}$ $\operatorname{\text{{vbll}}}$ 0.35 0.07 1. 0.31 1. 0.02 0.88 0.36 0.88 0.38
[…]
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+DiscVBLLMLP}}}$ $\operatorname{\text{{vbll}}}$ 0.85 0.08 0.97 0.27 0.89 0.08 0.95 0.27 0.94 0.35
[…]
Table 1: Comparison of estimators for the sentiment and factcheck datasets, with $\alpha$=0.95=0.95. R indicates all predictions were rejected, which is preferred over falling under the expected accuracy. $n=|\text{Admitted}|$, the count of non-rejected documents. The rows corresponding to the proposed $\operatorname{\text{{sdm}}}$ estimator (Eq. 9) are highlighted.
Co-variate-shifted and Out-of-distribution data.
With $\operatorname{\text{{SentimentOOD}}}$, the distinctions among the estimators become clear, with the non-$\operatorname{\text{{sdm}}}$-based estimators performing poorly, even in terms of marginal calibration. That would come as a surprise to end-users, whereas with the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ estimator, the out-of-distribution documents are more reliably rejected, with the few admitted predictions generally obtaining high conditional accuracies, despite the relatively low accuracies over the test set without selection (see $\operatorname{\text{{no-reject}}}$ in Table 2). A similar pattern is observed over the $\operatorname{\text{{Factcheck}}}$ dataset in Table 1, and Table 3 in the Appendix.
Understanding $q^{\prime}_{\rm{min}}$.
For reference, Table 4 provides the results over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for the $\operatorname{\text{{sdm}}}$-based estimators. The value of $q^{\prime}_{\rm{min}}$ tends to increase as the accuracy over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ decreases, reflecting a more conservative $\operatorname{\text{{high-reliability}}}$ region that admits fewer points. Alg. 1 failed to find a finite $q^{\prime}_{\rm{min}}$ for $\operatorname{\text{{Mixtral8x7B+sdm}}}$ over the $\operatorname{\text{{Factcheck}}}$ calibration set at $\alpha=0.95$, so for reference, we also show the $\operatorname{\text{{high-reliability}}}$ region at $\alpha=0.94$, as well as in Tables 1 and 3. In this way, $q^{\prime}_{\rm{min}}$ provides a principled, data-driven indicator of the reliability of the estimates, which is interpretable as a simple indicator as to whether the conditional accuracies are, or are not, obtainable over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ at the specified $\alpha$. Similar to conformal estimators, and unlike $\operatorname{\text{{tempScaling}}}$ and $\operatorname{\text{{vbll}}}$ estimators, Alg. 1 runs after the parameters of the adaptor layer have been fixed, so this indicator also in effect serves as a check on the optimization process (Eq. 3.1.6) itself.
Far out-of-distribution data.
The Appendix (Tables 5 and 6) provides results for the shuffled variants. The selection criteria of Eq. 9 reliably rejects the challenging predictions, whereas the non-$\operatorname{\text{{sdm}}}$-based estimators fare poorly, in general. In this way, the $\operatorname{\text{{sdm}}}$ activation serves as an effective out-of-distribution detection method. With existing methods, defining an out-of-distribution point has been task-specific, and generally challenging over high-dimensional inputs, typically requiring additional modeling beyond that of the calibration or selection method. In contrast, Eq. 9 provides a principled approach for determining such cut-offs in a data- and model-driven manner, with minimal hyper-parameters, resulting in a separation of points over which the estimator tends to be reliable (namely, the admitted points) and those over which the estimates themselves tend to be unreliable.
6 Conclusion
Limitations
The $\operatorname{\text{{sdm}}}$ estimator requires the $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ exemplar vectors at test-time, but the additional compute required for test-time matching is similar to that of commonly used dense retrieval mechanisms with LMs, so the additional overhead is likely achievable in typical use-cases.
The $\operatorname{\text{{sdm}}}$ activation function inherits the updatability property of instance-based metric learners. Instances with labels $y\in{\mathcal{Y}}=\{1,\ldots,C\}$ can be dynamically added to $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ after training. This can change the $\operatorname{\text{{Similarity}}}$ and $\operatorname{\text{{Distance}}}$ values, and by extension the uncertainty estimates, while the $\operatorname{\text{{Magnitude}}}$, and by extension the $\operatornamewithlimits{arg\,max}$ prediction, remain unchanged provided the weights of the CNN adaptor are held fixed. We hypothesize this can be a useful tradeoff between fast moving weights and slow moving weights in continual learning settings.
Relatedly, instances with labels $y\notin{\mathcal{Y}}$ can also be added to $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ after training. Given Eq. 3.1.2, test instances matching to such instances will have reduced $\operatorname{\text{$q$}}$ values, ceteris paribus, since the model can never predict such labels. This is potentially a useful, lightweight alternative to adding an explicit “out-of-distribution class” to $\mathcal{Y}$}, a comparison to which we leave to future work.
As indicated by our loss notation (Eq. 6) and described in § 3.1.2 and § 3.1.3, instance-wise $\operatorname{\text{$q$}}$ and $\operatorname{\text{$d$}}$ are not parameters learned directly via gradient descent. By design, we are not back-propagating through a bi- and/or cross-encoded search graph. The KNN approximations of Schmaltz (2021, §3.7.1) are trained with an iterative loss-masking approach to limit overfitting to outliers; with the $\operatorname{\text{{sdm}}}$ activation, instance-wise $\operatorname{\text{$q$}}$ and $\operatorname{\text{$d$}}$ serve as the regularizers during learning.
For clarity of presentation, in the main text we omit consideration of the Beta-distributed error term that is a function of the effective sample size of split-conformal coverage (Vovk, 2012). In practice, this term is negligible in the present experiments given $\lvert\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}\rvert$ and the resolution of the comparisons. See Appendix A.6 for a further discussion of analyzing the effective sample size for both the class-conditional and prediction-conditional estimates.
Among points in the $\operatorname{\text{{high-reliability}}}$ region, $\operatorname{\text{{sdm}}}$ estimators are relatively robust to covariate shifts that alter the relative proportion of points in the regions partitioned by the Similarity, Distance, and Magnitude signals, as well as label shifts, the latter due to the class-wise thresholds of Alg. 1. For example, in the extreme case, the proportion of points for which $\operatorname{\text{$q$}}=0$ and/or $\operatorname{\text{$d$}}=0$ can be much larger in the held-out $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ compared to that seen in training without altering the calibration of the $\operatorname{\text{{high-reliability}}}$ region, ceteris paribus. Of course, it is not possible to maintain calibration over all possible distribution shifts. However, in contrast, the examined non-$\operatorname{\text{{sdm}}}$-based calibration methods work quite poorly in maintaining calibration in high-probability regions in the presence of even modest distribution shifts, since they marginalize over those partitions, and they also marginalize over the labels. That limits the applicability of such calibration methods in practice. Importantly, due to the dense matching and partitioning of the calibration set, $\operatorname{\text{{sdm}}}$ estimators provide interpretability-by-exemplar: A user can examine the applicable documents in training and those in the applicable partition of the calibration set for further analysis and labeling, as needed.
Appendix A Appendix
The prompts for the tasks appear in Appendix A.1. Expanded versions of the tables in the main text appear in Appendix A.2, and the reference results over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ appear in Appendix A.3. We provide the results for the far out-of-distribution (OOD) shuffled datasets in Appendix A.4. Additional implementation details are included in Appendix A.5. Appendix A.6 provides an approach for analyzing the effective sample size for both the class-conditional and prediction-conditional estimates. We close with Appendix A.7, which provides additional details and comparisons for the experiments with variational Bayesian last-layer estimators (Harrison et al., 2024).
A.1 Prompts
Sentiment.
For the sentiment datasets, we prompt the LMs for a binary classification (Yes or No) as follows:
Here is a movie review. <review> DOCUMENT </review> Is the sentiment of the movie review positive? Answer Yes if the sentiment is positive. Answer No if the sentiment is negative. Start your response with Yes or No.
We replace DOCUMENT with the corresponding text for each instance.
Factcheck.
Similarly, for the factcheck datasets, we prompt the LMs for a binary classification (Yes or No) as follows:
Here is a statement that may contain errors. <statement> DOCUMENT </statement> Is the statement true? Answer Yes if the statement is true. Answer No if the statement is false. Start your response with Yes or No.
As above, we replace DOCUMENT with the corresponding text for each instance.
A.2 Additional Rows for Table 1
Additional rows for Table 1 appear in Table 2 for the sentiment datasets and Table 3 for the factcheck datasets.
A.3 Results over the Calibration Set
Table 4 provides reference results over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ to illustrate the behavior of $q^{\prime}_{\rm{min}}$.
A.4 Far OOD Shuffled Datasets
As discussed in Section 5 in the main text, Table 5 shows results for the $\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$ datasets, and Table 6 shows results for the $\operatorname{\text{{FactcheckShuffled}}}$ datasets.
In the case of $\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$, the semantics of the original labels are maintained. This requires the models and estimators to attempt a sentiment classification over the bag-of-words input, or reject the classification. This represents the setting where an LM is given far out-of-distribution input, and additionally provides a control on test-set contamination of the underlying LMs, which due to the shuffling, are relatively unlikely to have seen all of the long, contiguous n-gram sequences from these documents in training or fine-tuning.
In the case of $\operatorname{\text{{FactcheckShuffled}}}$, since the task prompt seeks to classify errors, we set the ground-truth labels of the shuffled counterparts to $y=0$.
A.5 Additional Implementation Details
We mean center the input to $g$, the 1-D CNN of the $\operatorname{\text{{sdm}}}$ activation layer and the otherwise identical CNN adaptors of the baseline comparison estimators, via the mean and standard deviation over $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. In all experiments with adaptor layers, $M=1000$ and we use a mini-batch size of 50. We use the Adam optimizer (Kingma and Ba, 2017) (without weight decay) with a learning rate of $1\times 10^{-5}$ for training.
A.5.1 Implementation of the SDM Activation Function
A.5.2 Implementation of the Empirical CDF Function
A.6 Analyzing the Effective Sample Size
Assumption A.1.
$\mathbf{\hat{n}}$= $[\lvert\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}\rvert^{y_{1}}\cdot{\rm{eCDF}}^{y_{1}}_{\rm{ca}}(q^{\prime}),\ldots,$
$~\lvert\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}\rvert^{y_{C}}\cdot{\rm{eCDF}}^{y_{C}}_{\rm{ca}}(q^{\prime})]$ (10)
where $\lvert\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}\rvert^{y_{c}}$ is the count of calibration set points with true label $y=c$.
For the prediction-conditional estimates, assuming independent and identically distributed (i.i.d.) data, these sample size estimates can be used to construct a band around the empirical CDFs over $\operatorname{\text{$d_{\rm{nearest}}$}}$ (Eq. 3.1.3) using the sharp constant (Massart, 1990) of the distribution-free DKW inequality (Dvoretzky et al., 1956), with $\hat{n}_{\rm{min}}$ taken as the minimum among the estimated sample sizes across classes for the test instance:
$\hat{n}_{\rm{min}}$ $=\min\left[\hat{n}_{1},\ldots,\hat{n}_{C}\right]$ (12)
If $\hat{n}_{\rm{min}}=0$, our convention is to set $\epsilon=1$. We then construct the conservative lower and upper counterparts to the distance quantile of Eq. 3.1.3:
A.7 Comparison to Bayesian Last-layer Networks
In this section, we provide additional details for the comparisons to variational Bayesian last-layer neural networks ($\operatorname{\text{{vbll}}}$), a computationally efficient Bayesian approach (Harrison et al., 2024). The basic setup is similar to the Frequentist and empirically motivated approaches using CNN-based adaptors examined in the main text in that it involves training a small final-layer adaptor over the frozen parameters of the language model. However, in this case, the adaptor network is a multi-layer perceptron (MLP) combined with the $\operatorname{\text{{vbll}}}$ estimator. We consider both the discriminative and generative $\operatorname{\text{{vbll}}}$ estimators.
Models.
We follow the parameter choices and architecture of Harrison et al. (2024), and its associated code tutorial [^7] Available at https://github.com/VectorInstitute/vbll, with applicable changes to match the experimental settings of the other models and estimators. Specifically, the VBLL models consist of an input linear layer, 2 core linear layers, and a final output linear layer. The hidden dimension is set at 795, which yields a similar number of parameters as the $\operatorname{\text{{sdm}}}$ activation layers (approximately 6 million parameters for the Phi-3.5-mini-instruct adaptors and approximately 8 million parameters for the Mixtral 8x7B Instruct v0.1 adaptors). The input to the VBLL models is the same mean-centered embeddings used with the $\operatorname{\text{{sdm}}}$ activation layers and the baseline CNN adaptor layers of the main text. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with a weight decay of $1\times 10^{-4}$, which following the code tutorial, is not applied to the final layer. We use a learning rate of $1\times 10^{-5}$ for training, matching that used with the $\operatorname{\text{{sdm}}}$ activation layers. Following the code tutorial, we use gradient clipping with a max norm of 1, and we use the exponential linear unit (ELU) as the activation function (Clevert et al., 2016). We train for 200 epochs, choosing the epoch with the lowest held-out validation loss over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ as the chosen weights. We repeat this process for $J=10$ shuffles of the data (i.e., splits of $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$), choosing the model with the globally lowest overall held-out validation loss over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ as the final model.
For the discriminative models over Phi-3.5-mini-instruct and Mixtral 8x7B Instruct v0.1, we use the labels $\operatorname{\text{{phi3.5+DiscVBLLMLP}}}$ and $\operatorname{\text{{Mixtral8x7B+DiscVBLLMLP}}}$, respectively. Those are the models appearing in Table 1 in the main text. For the generative models over Phi-3.5-mini-instruct and Mixtral 8x7B Instruct v0.1, we use the labels $\operatorname{\text{{phi3.5+GenVBLLMLP}}}$ and $\operatorname{\text{{Mixtral8x7B+GenVBLLMLP}}}$, respectively. For all of the aforementioned models, we use a KL regularization weight set at $\frac{1}{|\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}|}$, as in the original paper. We also consider analogous models that increase the KL regularization weight by a multiplicative factor of 50. Increasing the KL regularization weight is suggested in the code tutorial as the “simplest and most effective method to control the scale of uncertainty” of $\operatorname{\text{{vbll}}}$ models. For these models with the larger KL regularization weight, we use the labels $\operatorname{\text{{phi3.5+DiscVBLLMLP}}_{\text{rw50}}}50$, $\operatorname{\text{{Mixtral8x7B+DiscVBLLMLP}}_{\text{rw50}}}50$, $\operatorname{\text{{phi3.5+GenVBLLMLP}}_{\text{rw50}}}50$, and $\operatorname{\text{{Mixtral8x7B+GenVBLLMLP}}_{\text{rw50}}}50$, respectively.
We use the label $\operatorname{\text{{vbll}}}$ for the estimator that thresholds the output of the variational Bayesian last-layer neural network at $\alpha$ for the predicted class, analogous to the $\operatorname{softmax}$ and $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ estimators. The $\operatorname{\text{{no-reject}}}$ estimator provides a reference point without any selection criteria applied (i.e., the standard marginal and class- and prediction-conditional accuracies over the given dataset).
Results.
The results for the sentiment datasets appear in Table 7, and the results for the factcheck datasets appear in Table 8. For reference, in all cases we also provide the results over the held-out calibration set, $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, which is the held-out split used to determine the final model weights, as noted above.
Comparing to Table 4, the accuracies over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for the $\operatorname{\text{{no-reject}}}$ estimators are similar for the $\operatorname{\text{{vbll}}}$ models and the models using $\operatorname{\text{{sdm}}}$ activation functions. The MLPs of the $\operatorname{\text{{vbll}}}$ models and the 1-D CNNs of the $\operatorname{\text{{sdm}}}$ activation layers have a similar number of parameters and are trained with the same maximum number of epochs and the same number of iterated shuffles of $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. The accuracies without selection for the $\operatorname{\text{{Sentiment}}}$ calibration set are already at least $\alpha$, and those for the $\operatorname{\text{{Factcheck}}}$ calibration set are below $\alpha$, but at least $\alpha-0.05$. As such, differences in calibration effectiveness of the respective methods are not directly attributable to substantively different baseline accuracies for the in-distribution tasks.
We find that the $\operatorname{\text{{vbll}}}$ estimators are well-calibrated in high-probability regions over in-distribution data, but generally fare poorly over co-variate shifts and out-of-distribution data. We find no clear advantages nor disadvantages for the discriminative vs. generative variants, and modifying the KL regularization weight has a minimal impact, at least at this scale. In Harrison et al. (2024), $\operatorname{\text{{vbll}}}$ estimators over LMs for sentiment classification are only compared to an MLP baseline on in-distribution data; our evaluation setting is significantly more challenging and closer to real-world conditions encountered with LM applications.
@@ Results.
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 0.99 0.43 0.99 0.38 0.99 0.43 0.99 0.38 0.99 0.81
@@
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 0.98 0.48 0.98 0.43 0.98 0.47 0.98 0.43 0.98 0.90
@@
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 1. 0.01 0.98 0.01 0.98 0.01 1. 0.01 0.99 0.02
@@
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 0.98 0.05 0.96 0.04 0.97 0.05 0.98 0.04 0.97 0.08
@@
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 1. 0.01 0.97 0.14 0.75 0.02 1. 0.14 0.97 0.16
@@
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 0.88 0.10 0.95 0.18 0.91 0.09 0.93 0.18 0.93 0.27
@@
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}},\alpha=0.94$ 0.85 0.11 0.95 0.18 0.92 0.10 0.91 0.19 0.91 0.29
@@
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 0.99 0.42 0.99 0.39 0.99 0.42 0.99 0.39 0.99 0.81
@@
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 0.99 0.47 0.98 0.43 0.98 0.47 0.98 0.43 0.98 0.90
@@
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 0.98 0.33 0.99 0.27 0.99 0.33 0.98 0.27 0.98 0.60
@@
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ 0.97 0.41 0.95 0.32 0.96 0.41 0.95 0.32 0.96 0.73
@@
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\text{$\alpha$}},\alpha=0.94$ 0.96 0.42 0.95 0.33 0.96 0.42 0.95 0.32 0.96 0.74